OpenStudy (anonymous):

If 2^x=√2, te solution of (2^x +1)^2 = (4^x) +3 is...

OpenStudy (anonymous):

First thing's first, can you rewrite 4^x?

OpenStudy (anonymous):

Mmm... i don't understand....mm four to the x power?

OpenStudy (anonymous):

You ideally want to be able to get 4^x into a form where knowing that 2^x = √2 becomes helpful. Here's a hint: \[(a^b)^c = (a^c)^b\]

OpenStudy (anonymous):

X=1/2?

OpenStudy (anonymous):